A body execute SHM of time period 6s. If it's mass be 0.1kg, it's velocity, 1s after it passes mean position is 3m/s. Find its


A. Kinetic energy


B. Potential energy


C. Total energy


1
Expert's answer
2021-03-29T09:02:20-0400

a)


"K=0.5mv^2\\K=0.5(0.1)(3)^2=0.45\ J"

b)


"v=A\omega \cos{(\omega t)}\\3=A\frac{2\pi}{6} \cos{\frac{2\pi}{6} }\\x=\frac{18}{\pi}\ m"

"TE=0.5m\omega^2A^2\\TE=0.5(0.1)\frac{4\pi^2(18)^2}{\pi^26^2}=1.8\ J"

c)


"U=1.8-0.45=1.35\ J"